OpenStudy (anonymous):

Which of the points is a solution to 2x – 4y = 6? Somebody please help me! . A. (-2, 1) B. (1, -1) C. (1, 1) D. (2, 1)

OpenStudy (dan815):

plug in your point and see if it satisfies that equation

563blackghost (563blackghost):

\(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\sf\color{blue}{~~~~Welcome~to~OpenStudy!~\ddot\smile}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

jimthompson5910 (jim_thompson5910):

You forgot to post the list of points. But let's say that (0,2) was one of the points. So (x,y) = (0,2) meaning that x = 0 and y = 2 Plug x = 0 and y = 2 into the equation to get... 2x – 4y = 6 2*0 – 4*2 = 6 ... replace x with 0 and y with 2 0 - 8 = 6 -8 = 6 .. this equation is false since the last equation is false, this means (0,2) is NOT a solution to 2x – 4y = 6

jimthompson5910 (jim_thompson5910):

now that I look back up to your post, you did update to include the possible answers. So check each answer choice in the same way I checked (0,2)

jimthompson5910 (jim_thompson5910):

Let me know if you still need help or not